OpenStudy (anonymous):

the value of a uncirculated Mint State-65' 1950 Jefferson nickel minted in Denver is 3/2 the value of the 1945 nickel minted in Philadelphia in similar condition. Together the total value of the two nickels is $20. What is the value of each coin?

OpenStudy (anonymous):

D + P = 20 D = 3/2P

OpenStudy (anonymous):

sub D in to the top equality (3/2P) + P = 20 5/2P = 20 P = 20*(2/5) P = 8

OpenStudy (anonymous):

then put P back into either equation D + 8 = 20 D = 20 - 8 D = 12

OpenStudy (anonymous):

verify that (3/2)*8 = 12

OpenStudy (anonymous):

thanks that was bugger to figure out i have another